# Large Language Models in Argument Mining: A Survey

### v5

@@ 5 Argument Quality Assessment
Figure 3: Overview of quality dimensions for argument mining quality assessment discovered in the surveyed literature. Note: The figure is an extension of a taxonomy proposed in Ivanova et al. (2024)
Argument Quality Assessment (AQA)—also termed Argument Evaluation or Argumentation Quality (AQ), is the stage of the argument-mining pipeline that quantifies how well an already-identified argument fulfils its inferential and communicative objectives. Whereas argument component identification and relation detection answer the structural questions “what is an argument and how is it built?”, AQA addresses the normative question “how good is this argument?”. Early philosophical (Wenzel, 1990) syntheses distinguish three overarching dimensions, Logic/Cogency, Dialectic/Reasonableness, and Rhetoric/Effectiveness, each further decomposed into sub-dimensions such as acceptability, sufficiency, relevance, audience adaptation, or stylistic appropriateness (shown in Figure 3) .
@@
[…]
@@
Logic Acceptability Premises (and claim) do not contradict widely accepted facts or domain conventions (Wachsmuth et al., 2017a).
Local relevance Each premise bears a direct inferential link to its claim; off-topic or redundant information lowers the score (Wachsmuth et al., 2017b).
Sufficiency The collective evidence warrants the claim; absence of missing warrants yields a “fully supported” label (Stab and Gurevych, 2017; Gurcke et al., 2021).
Dialectic Global relevance The argument moves the overall discussion toward issue resolution—e.g., high in IBM Rank-30k “would-you-use” labels (Toledo et al., 2019; Gretz et al., 2020).
Opponent sensitivity Anticipates or rebuts likely objections, showing dialogical awareness (Khatib et al., 2017; Habernal and Gurevych, 2017).
Overall reasonableness Conforms to norms of critical discussion; free of straw man, ad hominem, and similar fallacies (Baff et al., 2018).
Rhetoric Credibility Projected expertise, trustworthiness, and goodwill via source citation, hedging, or professional register (Khatib et al., 2016).
Appropriateness Conformity of tone, politeness, and word choice to social norms and platform guidelines; penalizes toxicity or profanity (Ziegenbein et al., 2023).
Clarity Absence of ambiguity and needless complexity; logical flow and unambiguous wording facilitate comprehension (Persing et al., 2010).
Emotional Appeal Strategic use of sentiment, moral framing, or arousal to motivate the audience (Chen and Eger, 2025).
Arrangement Coherent discourse structure and persuasive ordering of premises, often assessed via discourse markers or paragraphing cues (Persing et al., 2010).
Strength Perceived inherent force or weight of the reasoning, independent of actual attitude change (Lukin et al., 2017).
Convincingness Likelihood that a neutral judge would find the argument compelling in pairwise or listwise comparisons (Gienapp et al., 2020).
Persuasiveness Observed ability to shift audience stance or votes after exposure; measured on platforms such as Change My View (Tan et al., 2016).
Winning Side Debate-level outcome: side whose cumulative arguments secure more post-debate votes or attitude change (Ivanova et al., 2024).
@@ 5.1 Quality Dimensions and Theoretical Frameworks
Argument quality is intrinsically multidimensional. During the last decade, two complementary lines of work have shaped the field: (i) theoretical systematisation, rooted in philosophy and argumentation studies (Wachsmuth and Werner, 2020); and (ii) empirical operationalisation, driven by the need to annotate large corpora and to train machine-learning models. We synthesise both strands in three steps.
@@ 5.1.1 Classical Tripartite Perspectives: Logic, Dialectic, Rhetoric
Wenzel’s “three perspectives” framework remains the most influential point of departure for computational AQ . It distinguishes a logical (cogency), a dialectical (reasonableness), and a rhetorical (effectiveness) vantage point on every argument (Wenzel, 1990). Wachsmuth et al. (2017a) operationalize the triad for NLP by posing three guiding questions:
@@
  • Effectiveness Does the discourse persuade its intended audience through credible, emotional, and well-structured presentation?
These perspectives provide the conceptual backbone for nearly all annotation schemes surveyed by Ivanova et al. (2024). The prevalence of the triptych is not merely historical; empirical work shows that annotators perceive and rate these facets with reasonably high reliability when clear instructions are provided (Gretz et al., 2020; Alhamzeh, 2023).
@@ 5.1.2 Fine-grained Sub-dimensions
As summarised in Table 3, each macro-perspective of the Logic–Dialectic–Rhetoric triad can be decomposed into finer, operational sub-dimensions. For the logical tier, empirical work most often labels acceptability, local relevance, and sufficiency. The dialectical tier adds discourse-level notions—global relevance, opponent sensitivity, and overall reasonableness. The rhetorical tier widens the lens still further, capturing audience-oriented aspects such as credibility, appropriateness, clarity, emotional appeal, arrangement, and high-level outcome variables (strength, convincingness, persuasiveness, winning side). Early studies showed that even non-experts agree substantially on some facets—most notably sufficiency (Stab and Gurevych, 2017). Yet boundaries remain fluid: “convincingness” blends logical sufficiency with emotional impact, while “persuasiveness” often collapses the full triad into a holistic score. Consequently, recent annotation protocols increasingly employ multi-label or soft-label designs to let a single argument express several quality aspects at once (Plank, 2022).
@@ 5.1.3 Comparative Taxonomy Analysis
Blair 2011 3 Philosophical foundation of the Logic Dialectic Rhetoric triad.
Wachsmuth et al. 2017a 15 Dagstuhl 15512 ArgQuality ($\approx$300 arguments) First computational taxonomy.
Lauscher et al. 2020 11 GAQCorpus ($\approx$5k arguments) Demonstrates large-scale crowd annotation.
Ivanova et al. 2024 15 Aligns overlapping labels across 211 papers; proposes harmonised quality meta-labels for transfer learning.
Nussbaum et al. 2020 7 98 science-teacher lesson plans Maps Walton’s Critical Questions to classroom feedback; emphasises opponent rebuttal and reasonableness.
Grundler et al. 2024 5 225 Italian legal decisions Adds statute relevance and argument-scheme fit; stresses domain-specific sufficiency and formality.
@@
As a example, Table 4 shows the changes of such theory framework, contrasts four influential taxonomies: (Blair, 2011), (Wachsmuth et al., 2017a), (Lauscher et al., 2020), (Ivanova et al., 2024) and two domain-specific schemes (legal AQ (Grundler et al., 2024), educational feedback (Nussbaum et al., 2020)). Three trends emerge: (i) Convergence on the tripartite backbone—every scheme retains some form of Logic–Dialectic–Rhetoric. (ii) Progressive granularity—the average number of annotated sub-dimensions grows from 3 (2010) to 11 (2024). (iii) Domain hybridisation new schemes borrow constructs from adjacent fields (fact-checking, toxicity detection, essay scoring), leading to overlapping labels that challenge cross-dataset transfer.
@@ 5.2 Modelling Advances: From Feature Engineering to LLM-based Judges
Early automatic approaches to argument-quality assessment relied on feature-engineering pipelines. Linear, SVM and gradient-boosting models consumed handcrafted indicators such as length, sentiment, hedge density and discourse connectives, delivering the first reproducible baselines on essays, reviews and Web debates (Habernal and Gurevych, 2017; Wachsmuth et al., 2017a). On Reddit Change My View, simple length and interaction features already yielded reliable convincingness rankings (Chalaguine and Schulz, 2017). Potash et al. were first to cast convincingness as a pairwise ranking problem, introducing a margin-based loss that remains standard today (Potash et al., 2019). Other work encoded topical similarity to improve local relevance (Gu et al., 2018), applied PageRank over argument graphs to estimate global salience (Wachsmuth et al., 2017b) or factual-consistency rules to flag insufficient support (Stab and Gurevych, 2017). Although feature lists were task-specific, they established the feasibility of automatic AQA on modest corpora such as AAE and Webis ArgQuality (Wachsmuth et al., 2017a).
@@
The 2016–2021 period marks an unequivocal shift from sparse inputs to representation learning. Bidirectional LSTMs outperformed bag-of-words on convincingness by capturing sequential cues (Habernal and Gurevych, 2016); CNN–BiLSTM hybrids added early cross-topic generalization on comment threads (Huber et al., 2019). Several groups then moved from classification to ranking losses: margin-based pairwise objectives (Potash et al., 2019) and listwise NDCG targets (Gu et al., 2018) yielded more stable quality orderings. Transformer fine-tuning largely removed feature design: BERT and RoBERTa lifts of 5–10 F1 or $\rho$ on IBM-Rank and GAQ tasks are now well documented (Gretz et al., 2020; Lauscher et al., 2020). Downstream improvements came from metric-learning variants—e.g. Siamese-SBERT with contrastive loss for out-of-domain convincingness and from multi-task heads that jointly predict stance or component type, boosting macro-F1 on low-resource sets (Favreau et al., 2022). Efficient pairwise crowdsourcing accelerated data creation and ranking-loss training (Gienapp et al., 2020), while contextual embeddings enabled intrinsic, reference-free scoring of short arguments (Wachsmuth and Werner, 2020). Together, these advances consolidated neural encoders as the default for AQA.
@@
Sentence-level encoders struggle with logical sufficiency and discourse flow, prompting richer context integration. Gurcke et al. generate candidate conclusions and use textual entailment to operationalise sufficiency (Gurcke et al., 2021). External-knowledge infusion—via retrieval-augmented transformers—raises sufficiency F1 by 3 points on legal AMELIA when factual statutes are injected at inference time (Liu et al., 2024b). Graph neural networks layered on RST or AMT structures improve global relevance, rebuttal recognition, and “winning-side” prediction, notably in legal and scientific debates (Zhang et al., 2023; Galassi et al., 2023). Semantic-network-based rankers outperform sequential baselines on cross-topic challenge sets by explicitly modelling premise-claim-evidence ties (Ruiz-Dolz et al., 2023). Finally, discourse-intrinsic objectives such as the COVERAGE metric capture informativeness without labels and serve as pre-training tasks that transfer to sufficiency and clarity (Khosravani et al., 2024). These structure- and knowledge-aware approaches close much of the gap between surface-form language models and the normative criteria demanded by argumentation theory.
@@
[…]
@@
Large language models have recently moved from feature generators to stand-alone adjudicators. With as few as three in-context exemplars, GPT-4 matches or exceeds fine-tuned RoBERTa on GAQCorpus for pairwise convincingness and sufficiency (Chen et al., 2024). Performance improves further when prompts elicit chain-of-thought rationales and apply self-consistency voting: Wang et al. report a 7-point macro-F1 gain on IBM-Rank after majority-voting across five diverse reasoning traces (Altemeyer et al., 2025). Retrieval-augmented prompting injects domain facts at inference time; on the AMELIA legal benchmark, RAG-GPT-4 lifts sufficiency F1 by an additional 3 points compared with vanilla GPT-4 (Liu et al., 2024b). Beyond accuracy, LLM judges reveal latent biases. Manipulation-check experiments demonstrate that GPT-4 systematically discounts emotionally charged but fallacious premises, whereas earlier transformer baselines over-score them (Chen and Eger, 2025). However, agreement may be inflated when both training data and evaluation sets share topic distributions (Schiller et al., 2022); hybrid pipelines therefore combine frozen LLM judges with lightweight re-rankers or calibration layers to reduce domain drift and latency (Ruosch et al., 2024; Kotelnikov et al., 2022).

### v6

@@ 5 Argument Quality Assessment
Figure 4: Overview of quality dimensions for argument mining quality assessment discovered in the surveyed literature. Note: The figure is an extension of a taxonomy proposed in DBLP:conf/emnlp/IvanovaHN24
Argument Quality Assessment (AQA)—also termed Argument Evaluation or Argumentation Quality (AQ), is the stage of the argument-mining pipeline that quantifies how well an already-identified argument fulfils its inferential and communicative objectives. Whereas argument component identification and relation detection answer the structural questions “what is an argument and how is it built?”, AQA addresses the normative question “how good is this argument?”. Early philosophical (wenzel1990three) syntheses distinguish three overarching dimensions, Logic/Cogency, Dialectic/Reasonableness, and Rhetoric/Effectiveness, each further decomposed into sub-dimensions such as acceptability, sufficiency, relevance, audience adaptation, or stylistic appropriateness (shown in Figure 4) .
@@
Formally, given a textual unit $a$, ranging from a single premise to a full argumentative thread, and, optionally, contextual information $c$ (topic, dialogue history, background knowledge), AQA seeks to output either (i) a scalar or categorical quality score $q(a,c)\in\mathbb{R}$, (ii) a vector of dimension-specific scores $\mathbf{q}(a,c)$, or (iii) a preference relation $a\succ b$ between pairs $(a,b)$. Approaches thus fall along two orthogonal axes: absolute vs. relative scoring—Likert-style ratings versus pairwise preferences, and intrinsic vs. extrinsic evaluation—judging the argument in isolation versus with external facts or opponent turns . Each design choice trades annotation complexity for construct validity: relative comparisons often raise inter-annotator agreement, yet absolute scales enable downstream aggregation and benchmarking.
@@
Logic Acceptability Premises (and claim) do not contradict widely accepted facts or domain conventions (DBLP:conf/eacl/WachsmuthSHPBHN17).
Local relevance Each premise bears a direct inferential link to its claim; off-topic or redundant information lowers the score (DBLP:conf/eacl/WachsmuthSA17).
Sufficiency The collective evidence warrants the claim; absence of missing warrants yields a “fully supported” label (DBLP:conf/eacl/GurevychS17; DBLP:conf/argmining/GurckeAW21).
Dialectic Global relevance The argument moves the overall discussion toward issue resolution—e.g., high in IBM Rank-30k “would-you-use” labels (DBLP:conf/emnlp/ToledoGCFVLJAS19; DBLP:conf/aaai/GretzFCTLAS20).
Opponent sensitivity Anticipates or rebuts likely objections, showing dialogical awareness (DBLP:conf/emnlp/KhatibWHS17; DBLP:journals/coling/HabernalG17).
Overall reasonableness Conforms to norms of critical discussion; free of straw man, ad hominem, and similar fallacies (DBLP:conf/conll/BaffWKS18).
Rhetoric Credibility Projected expertise, trustworthiness, and goodwill via source citation, hedging, or professional register (DBLP:conf/coling/KhatibWKHS16).
Appropriateness Conformity of tone, politeness, and word choice to social norms and platform guidelines; penalizes toxicity or profanity (DBLP:conf/acl/ZiegenbeinSLPW23).
Clarity Absence of ambiguity and needless complexity; logical flow and unambiguous wording facilitate comprehension (DBLP:conf/emnlp/PersingDN10).
Emotional Appeal Strategic use of sentiment, moral framing, or arousal to motivate the audience (DBLP:journals/corr/abs-2503-00024).
Arrangement Coherent discourse structure and persuasive ordering of premises, often assessed via discourse markers or paragraphing cues (DBLP:conf/emnlp/PersingDN10).
Strength Perceived inherent force or weight of the reasoning, independent of actual attitude change (DBLP:conf/eacl/WalkerALW17).
Convincingness Likelihood that a neutral judge would find the argument compelling in pairwise or listwise comparisons (DBLP:conf/acl/GienappSHP20).
Persuasiveness Observed ability to shift audience stance or votes after exposure; measured on platforms such as Change My View (DBLP:conf/www/TanNDL16).
Winning Side Debate-level outcome: side whose cumulative arguments secure more post-debate votes or attitude change (DBLP:conf/emnlp/IvanovaHN24).
@@ 5.1 Quality Dimensions and Theoretical Frameworks
Argument quality is intrinsically multidimensional. During the last decade, two complementary lines of work have shaped the field: (i) theoretical systematisation, rooted in philosophy and argumentation studies (DBLP:conf/coling/WachsmuthW20); and (ii) empirical operationalisation, driven by the need to annotate large corpora and to train machine-learning models. We synthesise both strands in three steps.
@@ 5.1.1 Classical Tripartite Perspectives: Logic, Dialectic, Rhetoric
Wenzel’s “three perspectives” framework remains the most influential point of departure for computational AQ . It distinguishes a logical (cogency), a dialectical (reasonableness), and a rhetorical (effectiveness) vantage point on every argument (wenzel1990three). DBLP:conf/eacl/WachsmuthSHPBHN17 operationalize the triad for NLP by posing three guiding questions:
@@
  • Effectiveness Does the discourse persuade its intended audience through credible, emotional, and well‐structured presentation?
These perspectives provide the conceptual backbone for nearly all annotation schemes surveyed by DBLP:conf/emnlp/IvanovaHN24. The prevalence of the triptych is not merely historical; empirical work shows that annotators perceive and rate these facets with reasonably high reliability when clear instructions are provided (DBLP:conf/aaai/GretzFCTLAS20; DBLP:phd/dnb/Alhamzeh23).
@@ 5.1.2 Fine-grained Sub-dimensions
As summarised in Table 3, each macro-perspective of the Logic–Dialectic–Rhetoric triad can be decomposed into finer, operational sub-dimensions. For the logical tier, empirical work most often labels acceptability, local relevance, and sufficiency. The dialectical tier adds discourse-level notions—global relevance, opponent sensitivity, and overall reasonableness. The rhetorical tier widens the lens still further, capturing audience-oriented aspects such as credibility, appropriateness, clarity, emotional appeal, arrangement, and high-level outcome variables (strength, convincingness, persuasiveness, winning side). Early studies showed that even non-experts agree substantially on some facets—most notably sufficiency (DBLP:conf/eacl/GurevychS17). Yet boundaries remain fluid: “convincingness” blends logical sufficiency with emotional impact, while “persuasiveness” often collapses the full triad into a holistic score. Consequently, recent annotation protocols increasingly employ multi-label or soft-label designs to let a single argument express several quality aspects at once (DBLP:conf/emnlp/Plank22).
@@ 5.1.3 Comparative Taxonomy Analysis
blair2011relationships blair2011relationships 3 Philosophical foundation of the Logic Dialectic Rhetoric triad.
DBLP:conf/eacl/WachsmuthSHPBHN17 DBLP:conf/eacl/WachsmuthSHPBHN17 15 Dagstuhl 15512 ArgQuality ($\approx$300 arguments) First computational taxonomy.
DBLP:conf/coling/LauscherNNT20 DBLP:conf/coling/LauscherNNT20 11 GAQCorpus ($\approx$5k arguments) Demonstrates large-scale crowd annotation.
DBLP:conf/emnlp/IvanovaHN24 DBLP:conf/emnlp/IvanovaHN24 15 Aligns overlapping labels across 211 papers; proposes harmonised quality meta-labels for transfer learning.
nussbaum2020using nussbaum2020using 7 98 science-teacher lesson plans Maps Walton’s Critical Questions to classroom feedback; emphasises opponent rebuttal and reasonableness.
DBLP:conf/clic-it/GrundlerGSFGPLS24 DBLP:conf/clic-it/GrundlerGSFGPLS24 5 225 Italian legal decisions Adds statute relevance and argument-scheme fit; stresses domain-specific sufficiency and formality.
@@
As a example, Table 4 shows the changes of such theory framework, contrasts four influential taxonomies: (blair2011relationships), (DBLP:conf/eacl/WachsmuthSHPBHN17), (DBLP:conf/coling/LauscherNNT20), (DBLP:conf/emnlp/IvanovaHN24) and two domain-specific schemes (legal AQ (DBLP:conf/clic-it/GrundlerGSFGPLS24), educational feedback (nussbaum2020using)). Three trends emerge: (i) Convergence on the tripartite backbone—every scheme retains some form of Logic–Dialectic–Rhetoric. (ii) Progressive granularity—the average number of annotated sub-dimensions grows from 3 (2010) to 11 (2024). (iii) Domain hybridisation new schemes borrow constructs from adjacent fields (fact-checking, toxicity detection, essay scoring), leading to overlapping labels that challenge cross-dataset transfer.
@@ 5.2 Modelling Advances: From Feature Engineering to LLM-based Judges
Early automatic approaches to argument-quality assessment relied on feature-engineering pipelines. Linear, SVM and gradient-boosting models consumed handcrafted indicators such as length, sentiment, hedge density and discourse connectives, delivering the first reproducible baselines on essays, reviews and Web debates (DBLP:journals/coling/HabernalG17; DBLP:conf/eacl/WachsmuthSHPBHN17). On Reddit Change My View, simple length and interaction features already yielded reliable convincingness rankings (DBLP:conf/eacl/ChalaguineS17). Potash et al. were first to cast convincingness as a pairwise ranking problem, introducing a margin-based loss that remains standard today (DBLP:conf/argmining/PotashFH19). Other work encoded topical similarity to improve local relevance (DBLP:conf/argmining/GuWXFLH18), applied PageRank over argument graphs to estimate global salience (DBLP:conf/eacl/WachsmuthSA17) or factual-consistency rules to flag insufficient support (DBLP:conf/eacl/GurevychS17). Although feature lists were task-specific, they established the feasibility of automatic AQA on modest corpora such as AAE and Webis ArgQuality (DBLP:conf/eacl/WachsmuthSHPBHN17).
@@
The 2016–2021 period marks an unequivocal shift from sparse inputs to representation learning. Bidirectional LSTMs outperformed bag-of-words on convincingness by capturing sequential cues (DBLP:conf/acl/HabernalG16); CNN–BiLSTM hybrids added early cross-topic generalization on comment threads (DBLP:conf/argmining/HuberTRDB19). Several groups then moved from classification to ranking losses: margin-based pairwise objectives (DBLP:conf/argmining/PotashFH19) and listwise NDCG targets (DBLP:conf/argmining/GuWXFLH18) yielded more stable quality orderings. Transformer fine-tuning largely removed feature design: BERT and RoBERTa lifts of 5–10 F1 or $\rho$ on IBM-Rank and GAQ tasks are now well documented (DBLP:conf/aaai/GretzFCTLAS20; DBLP:conf/coling/LauscherNNT20). Downstream improvements came from metric-learning variants—e.g. Siamese-SBERT with contrastive loss for out-of-domain convincingness and from multi-task heads that jointly predict stance or component type, boosting macro-F1 on low-resource sets (DBLP:conf/flairs/FavreauZB22). Efficient pairwise crowdsourcing accelerated data creation and ranking-loss training (DBLP:conf/acl/GienappSHP20), while contextual embeddings enabled intrinsic, reference-free scoring of short arguments (DBLP:conf/coling/WachsmuthW20). Together, these advances consolidated neural encoders as the default for AQA.
@@
Sentence-level encoders struggle with logical sufficiency and discourse flow, prompting richer context integration. Gurcke et al. generate candidate conclusions and use textual entailment to operationalise sufficiency (DBLP:conf/argmining/GurckeAW21). External-knowledge infusion—via retrieval-augmented transformers—raises sufficiency F1 by 3 points on legal AMELIA when factual statutes are injected at inference time (DBLP:conf/naacl/LiuFC24). Graph neural networks layered on RST or AMT structures improve global relevance, rebuttal recognition, and “winning-side” prediction, notably in legal and scientific debates (DBLP:conf/icail/ZhangNL23; DBLP:journals/taslp/GalassiLT23). Semantic-network-based rankers outperform sequential baselines on cross-topic challenge sets by explicitly modelling premise-claim-evidence ties (DBLP:conf/emnlp/Ruiz-DolzHG23). Finally, discourse-intrinsic objectives such as the COVERAGE metric capture informativeness without labels and serve as pre-training tasks that transfer to sufficiency and clarity (DBLP:conf/naacl/Khosravani0T24). These structure- and knowledge-aware approaches close much of the gap between surface‐form language models and the normative criteria demanded by argumentation theory.
@@
When labelled data are scarce, models can rely on intrinsic signals that approximate argumentative soundness without human scores. Wachsmuth & Werner’s Intrinsic QA predicts overall quality from discourse statistics—sentence length variance, connective density, argumentative role ratios—achieving $\rho=.46$ against crowdsourced judgments while using no external labels (DBLP:conf/coling/WachsmuthW20). Residual multi-task networks that jointly learn stance, component type, and quality raise macro-F1 on low-resource Italian legal arguments by 4–6 points over single-task baselines (DBLP:journals/taslp/GalassiLT23). Khosravani et al.’s COVERAGE metric uses masked-language-model perplexity to score how exhaustively a summary reflects key points, doubling as an unsupervised sufficiency proxy (DBLP:conf/naacl/Khosravani0T24). Finally, Plank’s variance analysis recommends treating disagreement as signal rather than noise, motivating the current shift toward soft or distributional quality labels (DBLP:conf/emnlp/Plank22).
@@
Large language models have recently moved from feature generators to stand-alone adjudicators. With as few as three in-context exemplars, GPT-4 matches or exceeds fine-tuned RoBERTa on GAQCorpus for pairwise convincingness and sufficiency (DBLP:conf/acl/ChenCLB24). Performance improves further when prompts elicit chain-of-thought rationales and apply self-consistency voting: Wang et al. report a 7-point macro-F1 gain on IBM-Rank after majority-voting across five diverse reasoning traces (DBLP:journals/corr/abs-2503-00847). Retrieval-augmented prompting injects domain facts at inference time; on the AMELIA legal benchmark, RAG-GPT-4 lifts sufficiency F1 by an additional 3 points compared with vanilla GPT-4 (DBLP:conf/naacl/LiuFC24). Beyond accuracy, LLM judges reveal latent biases. Manipulation-check experiments demonstrate that GPT-4 systematically discounts emotionally charged but fallacious premises, whereas earlier transformer baselines over-score them (DBLP:journals/corr/abs-2503-00024). However, agreement may be inflated when both training data and evaluation sets share topic distributions (DBLP:journals/corr/abs-2205-11472); hybrid pipelines therefore combine frozen LLM judges with lightweight re-rankers or calibration layers to reduce domain drift and latency (DBLP:conf/comma/RuoschLSB24; DBLP:journals/corr/abs-2206-09249).